# Hessian spectrum at the global minimum of the spherical pure-like mixed $p$-spin glasses

### v1

@@ Abstract
We study the large $N$-dimensional limit of the Hessian spectrum at the global minimum of a subclass of spherical mixed $p$-spin models, extending the relevant results in the regime of topology trivialization.
@@ 1 Introduction
Let $S^{N-1}(\sqrt{N})$ be the sphere of radius $\sqrt{N}$ in $N$-dimensional Euclidean space. The Hamiltonian of the spherical mixed $p$-spin model is given by
@@
Then direct calculation yields that the covariance of the Hamiltonian satisfies
@@
where $\Lambda_{N}$ is the normalized surface measure on the sphere $S^{N-1}(\sqrt{N})$. It is well-known that the limiting free energy can be computed through the famous Parisi formula [talagrand2006free], [chen2013aizenman]. From here, it can be deduced (cf. [AC17, (2) and (3)]) that the sequence $GS_{N}$ also converges to some value $GS$ almost surely, which is usually called limiting ground state energy,
@@
The Parisi formula for $GS$ through the variational representation of Parisi PDE was derived in [AC17, Theorem 1]. Meanwhile, based on the Crisanti-Sommers representation, the Parisi formula for the limiting ground state energy $GS$ was also established in [chen] and [2016dynamic].
@@
where the explicit expression of $\Theta_{p,k}(u)$ is given in [ABC13, (2.16)], which is commonly referred to as the complexity function of the pure $p$-spin model with index $k$. The function $\Theta_{p,k}(u)$ is known to be non-decreasing function. Let $-E_{0}(p)$ be the unique zero of function $\Theta_{p,0}(u)$, and define $E_{\infty}(p)=2\sqrt{\frac{p-1}{p}}$. Moreover, the second moment of $\operatorname{Crt}_{N}((-\infty,u))$ was calculated by [subag2017complexity], which showed that the critical points are concentrated around their means: for any $p\geq 3$ and $u\in\left(-E_{0}(p),-E_{\infty}(p)\right)$,
@@
This result was also obtained earlier in [ABC13, Theorem 2.12] for $p\geq 4$ even. A similar result [ABA13] was extended to the spherical mixed $p$-spin model, which states that for any open set $B\in\mathbb{R}$
@@
where $\theta_{k,\xi}(u)$ is explicitly given in [ABA13, (2.10)]. Setting $E_{\infty}^{\prime}(\xi)=\frac{2\xi^{\prime}(1)\sqrt{\xi^{\prime\prime}(1)}}{\xi^{\prime}(1)+\xi^{\prime\prime}(1)}$ as in [ABA13, (1.12)], the function $\theta_{k,\xi}(u)$ is strictly increasing on $\left(-\infty,-E_{\infty}^{\prime}(\xi)\right)$ and strictly decreasing on $\left(-E_{\infty}^{\prime}(\xi),\infty\right)$ with the unique maximum $\theta_{k,\xi}(-E_{\infty}^{\prime}(\xi))>0$. Denote by $-E_{0}(\xi)$ the smallest zero of $\theta_{0,\xi}(u)$. In Section 4 of their work, they claimed that $\xi(x)$ belonging to the $pure$-$like$ mixture (cf. Definition 1.1) class is equivalent to $E_{0}(\xi)=f_{1}$, where $f_{1}$ can be interpreted as the zero-temperature limit of one-step replica symmetry breaking ($1$-RSB) Parisi functional [ABA13, Propsition 6]. Also through a second moment calculation, Ben Arous, Subag and Zeitouni [BSZ20] proved that $GS=-E_{0}(\xi)$ for the $pure$-$like$ mixture models under Assumption I.
Recently, [Be22] considered the spherical mixed $p$-spin model on the unit sphere $S^{N-1}$ with external field
@@
where $h\geq 0$ is the strength of the external field and $\mathbf{u}_{N}\in S^{N-1}$ is a deterministic sequence. When $h^{2}>\xi^{\prime\prime}(1)-\xi^{\prime}(1)$, they showed that the Hamiltonian $H_{N}^{h}(\bm{\sigma})$ has trivial geometry, meaning that only critical points of $H_{N}^{h}$ are a single maximum and a single minimum. In this setting, they also found the convergence of the largest eigenvalues of the Hessian at global maximum [Be22, Theorem 1.1]. Indeed, the similar argument can be used to prove the convergence of the empirical spectrum measures of the Hessian at global maximum, with the help of a large deviation principle (LDP) for the spectrum measures of GOE matrices [benarous1997large]. The similar result was extended to locally isotropic Gaussian random fields on $\mathbb{R}^{N}$ [XZ22] provided the value of external field is large, which is previously known in the physics literature [FLD18].
@@
Then a natural question arises: what happens in the case of a small external field. For the general spherical mixed $p$-spin models, this may be difficult to handle, since it is very challenging to find the unique minimizer of the Parisi functional for the limiting ground state energy [chen, Theorem 1], and therefore the value of $GS$. However, it is known that $GS=-E_{0}(p)$ [subag2017complexity, Appendix IV] (see also [chen, Propisition 3]) for the pure $p$-spin model without external fields ($h$=0) and $GS=-E_{0}(\xi)$ [BSZ20, Theorem 1.1] for the $pure$-$like$ mixture models also without external fields (under Assumption I). This means that the smallest zero of the complexity function exactly matches the corresponding $GS$. Together with the analysis of complexity function, we are able to study the behavior of the Hessian spectrum at global minimum $\bm{\sigma}^{*}$ in the thermodynamic limit $N\rightarrow\infty$ for these subclass models, which is the main focus of this paper. For more general models, whether this property still holds remains unknown and requires more detailed investigation in the future. Furthermore, the $full$ (cf. Definition 1.1 below) mixture models contradict this above property (see [ABA13, Corollary 4.1]), which may pose a significant obstacle to solving this problem. Let
Definition 1.1.
A mixed model with covariance function $\xi(x)$ is called $pure$-$like$ if $G\left(\xi^{\prime},\xi^{\prime\prime}\right)>0$, $critical$ if $G\left(\xi^{\prime},\xi^{\prime\prime}\right)=0$, or $full$ if $G\left(\xi^{\prime},\xi^{\prime\prime}\right)<0$.
@@
From the definition, one can verify that pure $p$-spin model belongs to the $pure$-$like$ mixture class for all $p\geq 3$. However, due to the trivial relation (2.4) in pure $p$-spin model, we treat the pure $p$-spin case separately from the $pure$-$like$ but not pure $p$-spin case. For the latter case, we require the following condition as stated in [BSZ20].
@@
The explicit expression of the function $\Psi_{\xi}^{0}(r)$ is defined in [BSZ20, (1.11)]. Since it is not the key ingredient of this paper, we do not provide it here.
$\mathbb{E}(\mathrm{GOE}_{N})_{ij}^{2}=\frac{1+\delta_{ij}}{N}.$
We denote by $\mathcal{P}(\mathbb{R})$ the space of probability measures on $\mathbb{R}$, equipped with the bounded Lipschitz metric for two probability measures $\mu,\nu\in\mathcal{P}(\mathbb{R})$
where $\|f\|_{\infty}$ and $\|f\|_{L}$ denote the $L^{\infty}$ norm and Lipschitz constant of the function $f$, respectively. The empirical spectrum measure of an $N\times N$ matrix $M$ is denoted by $L_{N}(M)=\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}(M)}$, where $\lambda_{i}(M)$ are the eigenvalues of $M$. We further use $\lambda_{\min}(M)$ to denote the smallest eigenvalue of matrix $M$. Let $\sigma_{c,r}$ denote the semicircle measure with center $c$ and radius $r$, which has the following density with respect to Lebesgue measure
In particular, we denote the standard semicircle measure by $\sigma_{\rm sc}=\sigma_{0,2}$. To save space, we sometimes also use the shorthand notation $\xi^{\prime}=\xi^{\prime}(1)$ and $\xi^{\prime}=\xi^{\prime\prime}(1)$.
We organize the rest of the paper as follows. We prove the convergence of the empirical spectrum measures and smallest eigenvalues of the Hessian at global minimum for the spherical pure $p$-spin model in Section 2. Section 3 is dedicated to the analysis of spherical $pure$-$like$ mixture models.
@@ 2 Pure $p$-spin model
It is more convenient to work with this type of Gaussian process on the unit sphere $S^{N-1}$. For any $\bm{\sigma}\in S^{N-1}$, we define
@@
We endow the unit sphere $S^{N-1}$ with the standard Riemannian structure, induced by the Euclidean Riemannian metric on $\mathbb{R}^{N}$. We fix an orthonormal frame field $E=\left(E_{i}\right)_{i=1}^{N-1}$ on $S^{N-1}$, meaning that $\left(E_{i}(\bm{\sigma})\right)_{i=1}^{N-1}$ is a local orthonormal basis of $T_{\bm{\sigma}}S^{N-1}$ for any $\bm{\sigma}\in S^{N-1}$. We then define the spherical gradient and Hessian of $h_{N}(\bm{\sigma})$ as
@@
Furthermore, let $\partial_{r}h_{N}(\bm{\sigma})$ denote the radial derivative of $h_{N}(\bm{\sigma})$ at $\bm{\sigma}\in S^{N-1}$, which can be understood as the directional derivative of $h_{N}(\bm{\sigma})$ in Euclidean space in the direction $\bm{\sigma}$. For the relationship between $h_{N}(\bm{\sigma})$ and its derivatives, we refer the readers to [Be22, Lemma 3.2].
@@
$\begin{split}&\mathbb{E}\operatorname{Crt}_{N}(B,D)\\ &=\omega_{N}\varphi_{\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})}(0)\mathbb{E}\left\{\left\lvert\operatorname{det}(\nabla_{\mathrm{sp}}^{2}h_{N}(\bm{\sigma}))\right\rvert 1_{\left\{\frac{1}{\sqrt{N}}h_{N}(\bm{\sigma})\in B,\frac{1}{\sqrt{N}}\partial_{r}h_{N}(\bm{\sigma})\in D\right\}}\middle|\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})=0\right\},\end{split}$ (2.3)
@@
$\begin{split}&\mathbb{E}\operatorname{Crt}_{N}(B,D)\\ &=C_{p,N}\int_{D}\exp\left\{-\frac{Ny^{2}}{2p^{2}}\right\}\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right|1_{E}\left(\frac{y}{p}\right)\right\}\mathrm{d}y,\end{split}$ (2.6)
@@
$\begin{split}\lim_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{E}\mathrm{Crt}_{N}(B,D)&=\sup_{y\in D,y/p\in B}\left[\frac{1}{2}\log(p-1)-\frac{(p-2)y^{2}}{4p^{2}(p-1)}+\Phi\left(\frac{y}{\sqrt{2p(p-1)}}\right)\right]\\ &=:\sup_{y\in D,y/p\in B}R(y),\end{split}$ (2.7)
[…]
@@
where $\Theta_{p}(y)$ is the complexity function of spherical pure $p$-spin model given in [ABA13, (2.15)]. Thus, $R(y)$ is non-decreasing and continuous on $\mathbb{R}$, just like $\Theta_{p}(y)$. On the other hand, the limiting ground state energy for pure $p$-spin model was computed by Chen and Sen in [chen, Proposition 3]. Together with equality (1.3), if $\xi(x)=x^{p}$ for $p\geq 3$, then
@@
Recall the bounded Lipschitz metric $d$ between two probability measures given in (1.4). We exhibit our result for the pure $p$-spin model.
Theorem 2.1 (Pure $p$-spin model).
where $c_{p}=pE_{0}(p)$ with $E_{0}(p)$ given by (2.9) and $r_{p}=2\sqrt{p(p-1)}$.
Remark 2.1.
@@
$\mathrel{\phantom{=}}\frac{2\delta}{3\sqrt{p(p-1)}}$
$<d\left(L_{N-1}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right),\sigma_{c_{p}^{\prime},2}\right)$
@@
which implies $d(L_{N-1}(\mathrm{GOE}_{N-1}),\sigma_{\rm sc})>\frac{\delta}{3\sqrt{p(p-1)}}=:\delta^{\prime}$ by (2). Together with the Cauchy–Schwarz inequality, we get for $N$ large enough
$\mathbb{E}\mathcal{A}_{N,p}^{\varepsilon,\delta}$
$\leq C_{p,N}\int_{-c_{p}-p\varepsilon}^{-c_{p}+p\varepsilon}\exp\left\{-\frac{Ny^{2}}{2p^{2}}\right\}$
@@
$\mathbb{E}\mathcal{B}_{N,p}^{\varepsilon,\delta}$
$\leq C_{p,N}\int_{-c_{p}-p\varepsilon}^{-c_{p}+p\varepsilon}\exp\left\{-\frac{Ny^{2}}{2p^{2}}\right\}$
@@
$\frac{1}{N}H_{N}(\bm{\sigma})\in(-E_{0}(p)-\varepsilon,-E_{0}(p)+\varepsilon),\quad L_{N-1}(\nabla^{2}_{\mathrm{sp}}H_{N}(\bm{\sigma}))\notin B(\sigma_{c_{p},r_{p}},\delta).$
From the relation (2.1), the above event is equivalent to the event that there exists at least one critical points $\bm{\sigma}\in S^{N-1}$ with
@@
which completes the proof of (2.12). The above argument also can be applied to (2.13), since by Markov’s inequality, Lemma 2.3 directly gives $\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{B}_{N,p}^{\varepsilon,\delta}\geq 1)=0$. ∎

### v2

@@ Abstract
We study the large $N$-dimensional limit of the Hessian spectrum at the global minimum of a subclass of spherical mixed $p$-spin models, extending those results in the regime of topology trivialization.
@@ 1 Introduction
The Hamiltonians of spherical spin glasses are Gaussian smooth functions defined on the sphere in $N$-dimensional Euclidean space. Although the covariance function of the Hamiltonian depends only on the angle between two points, its landscapes are known to be rich and complex. The mean number of critical points (or those with a given index) and topology of its level sets of the Hamiltonian have been rigorously investigated in [ABA13]. On the other hand, the energy landscape near its global maximum was also examined in [Auffinger2018energy]. In this work, we will be interested in the behavior of Hessian spectrum at the global minimum (similarly at the global maximum, by symmetry).
We now introduce the model. Let $S^{N-1}(\sqrt{N})$ be the sphere of radius $\sqrt{N}$ in the $N$-dimensional Euclidean space. The Hamiltonian of the mixed spherical $p$-spin model is given by
@@
A direct calculation yields that the covariance of the Hamiltonian satisfies
@@
where $\Lambda_{N}$ is the normalized surface measure on the sphere $S^{N-1}(\sqrt{N})$. It is well-known that the limiting free energy exists and can be computed through the famous Parisi formula [talagrand2006free, chen2013aizenman]. From here, it can be deduced that the sequence $GS_{N}$ also converges to some value $GS$ almost surely (cf. [AC17, (2) and (3)]), which is usually called limiting ground state energy,
@@
Based on the Crisanti-Sommers representation, a variational formula for the limiting ground state energy $GS$ was also established in [chen], [2016dynamic], which should be seen as the analogue of the Parisi formula at zero temperature.
@@
where the explicit expression of $\Theta_{p,k}(u)$ is given in [ABC13, (2.16)], which is commonly referred to as the complexity function of the pure $p$-spin model with index $k$. The function $\Theta_{p,k}(u)$ is known to be non-decreasing function. Let $-E_{0}(p)$ be the unique zero of function $\Theta_{p,0}(u)$, and define $E_{\infty}(p)=2\sqrt{\frac{p-1}{p}}$. Moreover, the second moment of $\operatorname{Crt}_{N}((-\infty,u))$ was calculated in [subag2017complexity], which showed that the critical points are concentrated around their means: for any $p\geq 3$ and $u\in\left(-E_{0}(p),-E_{\infty}(p)\right)$,
@@
This result was also obtained earlier in [ABC13, Theorem 2.12] for $p\geq 4$ even. The complexity function was also calculated for the mixed spherical $p$-spin model [ABA13], which states that for any open set $B\in\mathbb{R}$,
@@
where $\theta_{k,\xi}(u)$ is explicitly given in [ABA13, (2.10)]. Setting $E_{\infty}^{\prime}(\xi)=\frac{2\xi^{\prime}(1)\sqrt{\xi^{\prime\prime}(1)}}{\xi^{\prime}(1)+\xi^{\prime\prime}(1)}$ as in [ABA13, (1.12)], the function $\theta_{k,\xi}(u)$ is strictly increasing on $\left(-\infty,-E_{\infty}^{\prime}(\xi)\right)$ and strictly decreasing on $\left(-E_{\infty}^{\prime}(\xi),+\infty\right)$ with the unique maximum $\theta_{k,\xi}(-E_{\infty}^{\prime}(\xi))>0$. Let
A mixed model with covariance function $\xi(x)$ is called $pure$-$like$ if $G\left(\xi^{\prime},\xi^{\prime\prime}\right)>0$, $critical$ if $G\left(\xi^{\prime},\xi^{\prime\prime}\right)=0$, or $full$ if $G\left(\xi^{\prime},\xi^{\prime\prime}\right)<0$. Denote by $-E_{0}(\xi)$ the smallest zero of $\theta_{0,\xi}(u)$. In Section 4 of their work, they claimed that $\xi(x)$ belonging to the $pure$-$like$ mixture class is equivalent to $E_{0}(\xi)=f_{1}$, where $f_{1}$ can be interpreted as the zero-temperature limit of one-step replica symmetry breaking ($1$-RSB) Parisi functional [ABA13, Propsition 6]. Also through a second moment calculation, Ben Arous, Subag and Zeitouni [BSZ20] proved that $GS=-E_{0}(\xi)$ for the $pure$-$like$ mixture models under Assumption I.
Recently, [Be22] considered the mixed spherical $p$-spin model on the unit sphere $S^{N-1}$ with an external field
@@
where $h\geq 0$ is the strength of the external field and $\mathbf{u}_{N}\in S^{N-1}$ is a deterministic sequence. When $h^{2}>\xi^{\prime\prime}(1)-\xi^{\prime}(1)$, they showed that the Hamiltonian $H_{N}^{h}(\bm{\sigma})$ has trivial geometry, meaning that the only critical points of $H_{N}^{h}$ are a single maximum and a single minimum. It is clear that the property above implies the vanishing of the maximum value of the complexity function. In this setting, they also found the convergence of the largest eigenvalues of the Hessian at global maximum [Be22, Theorem 1.1]. In fact, similar arguments can be used to prove the convergence of the empirical spectrum measures of the Hessian at global maximum, with the help of a large deviation principle (LDP) for the spectrum measures of the GOE matrices [benarous1997large]. Later, [XZ22] also confirmed the trivial topology phenomenon of locally isotropic Gaussian random fields on $\mathbb{R}^{N}$ and studied the convergence of empirical spectrum measures of the Hessian at global minimum provided the value of external field is large, which is previously known in the physics literature [FLD18].
@@
Out of the topology trivialization regime, it may be difficult to find the limiting Hessian spectrum at global minimum, since the maximum value of complexity function is positive, so that with high probability, there are exponentially many critical points. However, it is known that $GS=-E_{0}(p)$ [subag2017complexity, Appendix IV] (see also [chen, Propisition 3]) for the pure $p$-spin model without external fields ($h$=0) and $GS=-E_{0}(\xi)$ [BSZ20, Theorem 1.1] for the $pure$-$like$ mixture models also without external fields (under Assumption I). This means that the smallest zero of the complexity function exactly matches the corresponding $GS$. Together with the analysis of the complexity function, we can study the behavior of the Hessian spectrum at global minimum $\bm{\sigma}^{*}$ in the thermodynamic limit $N\rightarrow\infty$ for these subclass models, which is the main focus of this paper. For more general models, whether this property still holds remains unknown and will require a more detailed investigation in the future. Furthermore, the $full$ mixture models contradict this property above (see [ABA13, Corollary 4.1]), which may pose a significant obstacle to solving this problem.
1.1 Notations and main results
$\mathbb{E}(\mathrm{GOE}_{N})_{ij}^{2}=\frac{1+\delta_{ij}}{N}.$
We denote by $\mathcal{P}(\mathbb{R})$ the space of probability measures on $\mathbb{R}$, equipped with the bounded Lipschitz metric for two probability measures $\mu,\nu\in\mathcal{P}(\mathbb{R})$,
where $\|f\|_{\infty}$ and $\|f\|_{L}$ denote the $L^{\infty}$ norm and Lipschitz constant of the function $f$, respectively. The empirical spectrum measure of an $N\times N$ matrix $M$ is denoted by $L_{N}(M)=\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}(M)}$, where $\lambda_{i}(M)$ are the eigenvalues of $M$. We also use $\lambda_{\min}(M)$ to denote the smallest eigenvalue of matrix $M$. Let $\sigma_{c,r}$ denote the semicircle measure with center $c$ and radius $r$, which has the following density with respect to Lebesgue measure
In particular, we denote the standard semicircle measure by $\sigma_{\rm sc}=\sigma_{0,2}$. To save space, we sometimes also use the shorthand notation $\xi^{\prime}=\xi^{\prime}(1)$ and $\xi^{\prime\prime}=\xi^{\prime\prime}(1)$.
Theorem 1.1 (Pure $p$-spin model).
where $c_{p}=pE_{0}(p)$ with $E_{0}(p)$ given by (2.9) and $r_{p}=2\sqrt{p(p-1)}$.
@@ 1.1 Notations and main results
One can easily verify that the pure $p$-spin model belongs to the $pure$-$like$ mixture class for all $p\geq 3$. However, the particular trivial relation (2.4) exists only in the pure $p$-spin model. Therefore, for the sake of clear presentation, we treat the pure $p$-spin case separately from the $pure$-$like$ but not pure $p$-spin case. For the latter case, we require the following condition as stated in [BSZ20].
@@
The explicit expression of the function $\Psi_{\xi}^{0}(r)$ is defined in [BSZ20, (1.11)]. Since it is not the key ingredient in this paper, we do not provide it here.
Theorem 1.2 (Pure-like mixture models).
where $y_{0}$ is defined in (3.7).
Remark 1.1.
Recently, one month after we first posted this article on arXiv, Sellek [sellke2024marginal] deduced the limiting Hessian spectrum for even spins (i.e., $\gamma_{p}=0$ for all $p$ odd) via a remarkable construction of the unique minimizer of the Parisi functional of vector $p$-spin model [Zhou2022]. In particular, we emphasize that both our results, Theorems 1.1 and 1.2, include the odd terms. Moreover, the techniques used in our work differ from those in [sellke2024marginal] and should be of independent interest.
We organize the remainder of the paper as follows. We prove the convergence of the empirical spectrum measures and smallest eigenvalues of the Hessian at global minimum for the spherical pure $p$-spin model in Section 2. Section 3 is dedicated to the analysis of spherical $pure$-$like$ mixture models.
@@ 2 Pure $p$-spin model
It is more convenient to work with this type of Gaussian process on the $(N-1)$-dimensional unit sphere $S^{N-1}$. For any $\bm{\sigma}\in S^{N-1}$, we define
@@
We endow the unit sphere $S^{N-1}$ with the standard Riemannian structure, induced by the Euclidean Riemannian metric on $\mathbb{R}^{N}$. We fix an orthonormal frame field $E=\left(E_{i}\right)_{i=1}^{N-1}$ on $S^{N-1}$, which means that $\left(E_{i}(\bm{\sigma})\right)_{i=1}^{N-1}$ is a local orthonormal basis of $T_{\bm{\sigma}}S^{N-1}$ for any $\bm{\sigma}\in S^{N-1}$. We then define the spherical gradient and Hessian of $h_{N}(\bm{\sigma})$ as
@@
Furthermore, let $\partial_{r}h_{N}(\bm{\sigma})$ denote the radial derivative of $h_{N}(\bm{\sigma})$ at $\bm{\sigma}\in S^{N-1}$, which can be understood as the directional derivative of $h_{N}(\bm{\sigma})$ in Euclidean space in the direction $\bm{\sigma}$. For the relationship between $h_{N}(\bm{\sigma})$ and its derivatives, we refer the reader to [Be22, Lemma 3.2].
@@
$\begin{split}\mathbb{E}\operatorname{Crt}_{N}(B,D)=\omega_{N}\varphi_{\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})}(0)\mathbb{E}\left\{\left\lvert\operatorname{det}(\nabla_{\mathrm{sp}}^{2}h_{N}(\bm{\sigma}))\right\rvert 1_{\left\{\frac{1}{\sqrt{N}}h_{N}(\bm{\sigma})\in B,\frac{1}{\sqrt{N}}\partial_{r}h_{N}(\bm{\sigma})\in D\right\}}\middle|\nabla_{\mathrm{sp}}h_{N}(\bm{\sigma})=0\right\},\end{split}$ (2.3)
@@
$\begin{split}\mathbb{E}\operatorname{Crt}_{N}(B,D)=C_{p,N}\int_{D}\exp\left\{-\frac{Ny^{2}}{2p^{2}}\right\}\mathbb{E}\left\{\left|\operatorname{det}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right)\right|1_{E}\left(\frac{y}{p}\right)\right\}\mathrm{d}y,\end{split}$ (2.6)
@@
$\begin{split}\lim_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{E}\mathrm{Crt}_{N}(B,D)&=\sup_{y\in D\cap pB}\left[\frac{1}{2}\log(p-1)-\frac{(p-2)y^{2}}{4p^{2}(p-1)}+\Phi\left(\frac{y}{\sqrt{2p(p-1)}}\right)\right]\\ &=:\sup_{y\in D\cap pB}R(y),\end{split}$ (2.7)
where $\Phi(x)=\left[-\frac{|x|\sqrt{x^{2}-2}}{2}+\ln\left(\frac{|x|+\sqrt{x^{2}-2}}{\sqrt{2}}\right)\right]1_{\{|x|\geq\sqrt{2}\}}.$ In fact, due to the relation (2.4), one can check that for any $y\in\mathbb{R}$,
@@
where $\Theta_{p}(y)$ is the complexity function of spherical pure $p$-spin model given in [ABA13, (2.15)]. Thus, $R(y)$ is non-decreasing and continuous on $\mathbb{R}$, similar to $\Theta_{p}(y)$. On the other hand, the limiting ground state energy for pure $p$-spin model was computed by Chen and Sen [chen, Proposition 3]. Together with equality (1.3), if $\xi(x)=x^{p}$ for $p\geq 3$, then
@@
$\mathrel{\phantom{=}}\frac{2\delta}{3\sqrt{p(p-1)}}$ $<d\left(L_{N-1}\left(\mathrm{GOE}_{N-1}-\frac{\sqrt{N}y}{\sqrt{(N-1)p(p-1)}}\mathbf{I}_{N-1}\right),\sigma_{c_{p}^{\prime},2}\right)$
@@
which implies $d(L_{N-1}(\mathrm{GOE}_{N-1}),\sigma_{\rm sc})>\frac{\delta}{3\sqrt{p(p-1)}}=:\delta^{\prime}$ by (2). Together with the Cauchy–Schwarz inequality, we find for $N$ large enough
$\mathbb{E}\mathcal{A}_{N,p}^{\varepsilon,\delta}$ $\leq C_{p,N}\int_{-c_{p}-p\varepsilon}^{-c_{p}+p\varepsilon}\exp\left\{-\frac{Ny^{2}}{2p^{2}}\right\}$
@@
$\mathbb{E}\mathcal{B}_{N,p}^{\varepsilon,\delta}$ $\leq C_{p,N}\int_{-c_{p}-p\varepsilon}^{-c_{p}+p\varepsilon}\exp\left\{-\frac{Ny^{2}}{2p^{2}}\right\}$
@@
$\frac{1}{N}H_{N}(\bm{\sigma})\in(-E_{0}(p)-\varepsilon,-E_{0}(p)+\varepsilon),\quad L_{N-1}(\nabla^{2}_{\mathrm{sp}}H_{N}(\bm{\sigma}))\notin B(\sigma_{c_{p},r_{p}},\delta).$ (2.24)
Thus, the above event (2.24) is equivalent to the event that there exists at least one critical points $\bm{\sigma}\in S^{N-1}$ with
@@
which completes the proof of (1.5). The argument above also can be applied to (1.6), since by Markov’s inequality, Lemma 2.2 directly gives $\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{B}_{N,p}^{\varepsilon,\delta}\geq 1)=0$. ∎